OpenStudy (rock_mit182):

this is tough one, is prob.. In one version of a popular lottery game, a player selects six of the numbers from 1 to 54. The agency in charge of the lottery also selects six numbers. What is the probability that the player will match the six numbers if two50¢ tickets are purchased? (This jackpot is worth at least $2 million in prize money and grows according to the number of tickets sold.)

OpenStudy (rock_mit182):

I'm not good in probability please help

myininaya (myininaya):

Can a number be choose twice from 1 to 54?

OpenStudy (rock_mit182):

i guess so

myininaya (myininaya):

so on a smaller example if I could choose 2 numbers from1 through 4 (and I could choose a number from 1 through 4 more than once) Some combinations could be: 11 12 13 14 21 22 23 24 31 32 33 34 41 42 43 44 That is how many combinations: 4(4)=16. Which makes since. Because there are 4 ways you can choose the first number and 4 ways to choose the second number. Now one of these 16 will match the combination chosen by the people who run the lottery. So the probability of your combination being selected should be 1/16 for this example. right?

OpenStudy (rock_mit182):

but i don't get it

OpenStudy (rock_mit182):

c(n,k) =\[\frac{ n! }{ k!(n-k)! }\]

OpenStudy (rock_mit182):

it wouldn't be : \[\frac{ 54....48! }{ 6!(48!) }\]

OpenStudy (rock_mit182):

?

myininaya (myininaya):

you said I could choose a number more than once right?

OpenStudy (rock_mit182):

im lost

OpenStudy (rock_mit182):

why is my aproach wrong ?

OpenStudy (rock_mit182):

ok yeah . .

myininaya (myininaya):

oops sorry 54*54*54*54*54*54 is what i meant

myininaya (myininaya):

54^6

myininaya (myininaya):

just like on the smaller example we did 4*4 because we could select 4 numbers both both picks

OpenStudy (rock_mit182):

but what you are saying is not a permutation with repetition?

myininaya (myininaya):

I'm going off what you told me.... You said we could choose a number more than once.

myininaya (myininaya):

So you could choose 54 54 54 54 54 54 or 54 54 54 53 or so on...

myininaya (myininaya):

54 54 54 54 54 53*

OpenStudy (rock_mit182):

take it easy. well isaid that because i assumed the question says that

OpenStudy (rock_mit182):

could just we get the point

myininaya (myininaya):

I'm not upset. I'm just telling you that you told me that we could repeat a number.

OpenStudy (rock_mit182):

well i guess you read the problem too

myininaya (myininaya):

I was asking to clarify the problem. I have never done the lottery thing so I don't know how the lottery works. Can we pick a number more than once for our six numbers?

myininaya (myininaya):

Does the order matter in which we pick them?

OpenStudy (rock_mit182):

i know some of the principles, i just don't know how use the tools

OpenStudy (rock_mit182):

well is possible but you know the chances are so low

myininaya (myininaya):

it is possible the order doesn't matter in your problem?

myininaya (myininaya):

I would have to read how the lottery works I guess.

OpenStudy (rock_mit182):

br b

OpenStudy (rock_mit182):

\[\Delta = \left[\begin{matrix}a & -b \\ a & b\end{matrix}\right]\]

OpenStudy (rock_mit182):

\[x=\left[\begin{matrix}-1 & -b\\ 7& b\end{matrix}\right]\]/ delta \[y =\left[\begin{matrix}a & r \\ a& 7\end{matrix}\right]/ \Delta\]

OpenStudy (rock_mit182):